# On the Price of Decentralization in Decentralized Detection

### v1

@@ I-A Contribution
In this work, the above questions are addressed in the case of binary detection. We propose a generalization of the social learning rule in [2] and characterize the error exponents using tools in large deviation theory [10]. As a result, the error exponents of the original learning rule in [2] are characterized, which turn out to be strictly smaller than the error exponents in the centralized case. The reason is that the decentralized sources are not weighted equally due to the convergence of the Markov chain governing the consensus. Figure 1 illustrates the gap in error exponents with a simple example. In the example, 300 scale-free networks with 100 nodes in each are sampled. Each node serves as an independent Bernoulli source having consensus weights uniformly distributed to its neighbors. Gathering the consensus weights into a right stochastic matrix, the Markov chain with such corresponding transition matrix induces a unique stationary distribution denoted by $\pi$ under some minor assumptions. The figure shows that the error exponent of the original learning rule decreases with the network imbalance. We quantify the imbalance of the network with the 2-norm between $\pi$ and the uniform stationary distribution with each entry being 0.01 for this case. Notice that only when the network is balanced, the original learning rule obtain the optimal error exponent depicted by the blue dashed line.
@@ I-C Paper Organization
The rest of this paper is organized as follows. In Section II, we formulate our problem and introduce the learning rule proposed in [2]. In Section III, we propose our modified learning rule and show our main results. The proofs are then provided in Appendices. We proposed alternative learning rules for estimating the needed parameters and discuss the convergence of the estimation in Section IV. In Section V, we provide simulation results on the impact of network imbalance, estimation, and quantization. We make several discussions in Section VI including removing the assumptions on the network and extending our results to the multiple hypothesis testing problems. Finally, we make a brief conclusion in Section VII.
@@ II-B Social Learning Rule
The coefficients $Z^{(t)}_{i,1},Z^{(t)}_{i,2}$ in steps 2 and 4 normalize the belief vectors such that they fall back into the $m$-dimensional probability simplex.
@@
Though [2] characterized the convergence performance of the belief vectors, they did not study the probability of error, which seems to be a more concerned perspective in the conventional hypothesis testing problem. We will soon show that the learning rule proposed in [2] suffers a gap in the error exponent compared to the centralized case in Section II-D. Before then, we introduce the probability of error we consider in the rest of our work.
II-C Log-Belief Ratio Test and the Probability of Error
For the centralized binary detection problem, the randomized likelihood ratio test is optimal (in the Neyman-Pearson problem and the Bayes setting). However, in the decentralized setting, none of the nodes knows the joint likelihood of all the observations in the network and thus no one can carry out the likelihood ratio test. Under the above-mentioned learning rule, we consider the binary hypothesis testing problem, and a natural test based on the private belief vector maintained by each node emerges, which is defined as follows.
@@ II-C Log-Belief Ratio Test and the Probability of Error
Next, let us define the two types of probabilities of error for the log-belief ratio test.
Definition 3 (Probability of Error).
The type-I and type-II error probabilities for each node $i$ denoted by $\alpha_{i}^{(t)}(\eta_{i}^{(t)},\gamma_{i}^{(t)})$ and $\beta_{i}^{(t)}(\eta_{i}^{(t)},\gamma_{i}^{(t)})$ are defined as
$\alpha_{i}^{(t)}(\eta_{i}^{(t)},\gamma_{i}^{(t)})\triangleq$ $\;\mathsf{Pr}\big{\{}\varphi_{i}^{(t)}(\ell_{i}^{(t)})=1\;\big{|}\;\mathcal{H}_{0}\big{\}},$
$\beta_{i}^{(t)}(\eta_{i}^{(t)},\gamma_{i}^{(t)})\triangleq$ $\;\mathsf{Pr}\big{\{}\varphi_{i}^{(t)}(\ell_{i}^{(t)})=0\;\big{|}\;\mathcal{H}_{1}\big{\}}.$
Note that the performance depends on the parameters $\eta_{i}^{(t)},\gamma_{i}^{(t)}$.
It is then straightforward to come up with a decentralized Neyman-Pearson problem:
$\beta_{i}^{(t)*}(\epsilon)\triangleq$ $\underset{\eta_{i}^{(t)},\gamma_{i}^{(t)}}{\mathsf{minimize}}\;\;$ $\;\beta_{i}^{(t)}(\eta_{i}^{(t)},\gamma_{i}^{(t)})$
$\mathsf{subject\,to}$ $\;\alpha_{i}^{(t)}(\eta_{i}^{(t)},\gamma_{i}^{(t)})<\epsilon$
for all $i\in[n]$ with some $\epsilon\in(0,1)$. Our goal is to investigate the asymptotic behavior of $\beta_{i}^{(t)*}(\epsilon)$, that is, $\lim_{t\rightarrow\infty}-\frac{1}{t}\log\beta_{i}^{(t)*}(\epsilon).$ For the Bayes setting with prior $(\xi_{0},\xi_{1})$, we would consider the asymptotic behavior of the Bayes risk:
$\mathsf{P}_{\mathsf{e},i}^{(t)*}(r)\triangleq\min_{\eta_{i}^{(t)},\gamma_{i}^{(t)}}\left\{\xi_{0}\alpha_{i}^{(t)}(\eta_{i}^{(t)},\gamma_{i}^{(t)})+\xi_{1}\beta_{i}^{(t)}(\eta_{i}^{(t)},\gamma_{i}^{(t)})\right\}.$
II-D The Error Exponent Gap
As a brief glimpse into why we would adopt modifications on the original learning rule, we show that there is a gap between the error exponent in the centralized case and the one under the learning rule introduced in [2] through Theorem 1 in Section III-B.
Recall the weight matrix $W$ in the learning rule in Section II-B. The stochastic matrix could be seen as the transition matrix of a Markov chain. Under making some assumptions on the Markov chain later in Section III-B, the stationary distribution is unique and let it denoted by $\pi\in\Delta_{n}$. We say that the network is imbalanced if the stationary distribution $\pi$ is far from the uniform one, $\pi_{\mathrm{uniform}}=\frac{1}{n}\boldsymbol{1}$.
In Figure 1, as we briefly discussed in Section I, it can be observed that if the network tends to be more imbalance, the learning rule in [2] tends to suffer a larger gap in the error exponent compared to the centralized case, where a central node is assumed to be capable of gathering all the observations and likelihood functions to perform an optimal test. The gap between the error exponents motivates our modified learning rule introduced in the next section. It turns out that we can close the gap with slight modifications while keeping our learning rule working in a decentralized manner.
@@ III-A Modified Learning Rule
Now that the additional parameters $r_{i}$’s emerge, the probability of error depends on the choice of the geometric weights. Thus, we re-write our definitions in Section II-C.
@@
$\mathsf{P}_{\mathsf{e},i}^{(t)*}(r;\xi)\triangleq\min_{\eta_{i}^{(t)},\gamma_{i}^{(t)}}\left\{\xi_{0}\alpha_{i}^{(t)}(r;\eta_{i}^{(t)},\gamma_{i}^{(t)})+\xi_{1}\beta_{i}^{(t)}(r;\eta_{i}^{(t)},\gamma_{i}^{(t)})\right\}.$
@@ III-B First-Order Results
For the Neyman-Pearson problem in Section II-C, the type-II error exponent for each node $i$ is characterized as follows: for all $\epsilon>0$, the error exponent is
$\lim_{t\rightarrow\infty}-\frac{1}{t}\log\beta_{i}^{(t)*}(r,\epsilon)=\sup_{\lambda\geq 0}\left\{\sum_{j=1}^{n}\lambda\pi_{j}r_{j}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]-\log\mathsf{E}_{P_{1}}\left[\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda\pi_{j}r_{j}}\right]\right\}.$
for all $i\in[n]$.
@@
If each agent $i$ knows $\pi_{i}$, the value of the stationary distribution of the weight matrix at that node, by choosing $r=r^{*}$ where $r^{*}_{i}=c/\pi_{i}$ for all $i$ and some common constant $c\in\mathbb{R}$ among the nodes, we have
$\lim_{t\rightarrow\infty}-\frac{1}{t}\log\beta_{i}^{(t)*}(r^{*},\epsilon)=\sum_{j=1}^{n}D_{\mathrm{KL}}\left(P_{j,0}\|P_{j,1}\right)=D_{\mathrm{KL}}\left(P_{0}\|P_{1}\right)$
for all $i\in[n]$. The notation $P_{\theta}$ denotes the product distribution of $P_{1,\theta},...,P_{n,\theta}$ and $D_{\mathrm{KL}}$ denotes the Kullback–Leibler divergence.
@@
If each agent $i$ knows $\pi_{i}$, the value of the stationary distribution of the weight matrix at that node, by choosing $r=r^{*}$, for all prior $\xi$, we have
@@
where $\mathrm{CI}(P_{0},P_{1})$ denote the Chernoff information between $P_{0}$ and $P_{1}$, that is,
$\mathrm{CI}(P_{0},P_{1})=\max_{\lambda\in[0,1]}\left\{-\log\mathsf{E}_{P_{0}}\left[\left(\frac{P_{1}(\boldsymbol{X})}{P_{0}(\boldsymbol{X})}\right)^{\lambda}\right]\right\}=\max_{\lambda\in[0,1]}\left\{-\log\mathsf{E}_{P_{0}}\left[\left(\frac{P_{1,1}(X_{1})\dots P_{n,1}(X_{n})}{P_{1,0}(X_{1})\dots P_{n,0}(X_{n})}\right)^{\lambda}\right]\right\}$
@@ III-C Higher-Order Asymptotics
While Theorem 2 shows that decentralization does not affect the error exponent if we choose the geometric weights properly, further investigation on the probability of error may reveal the effect of decentralization. To characterize the higher-order asymptotics, further information about the distributions is needed. Let $S_{n}^{2}$ and $\alpha_{n}$ denote the second and third central moment of $\sum_{i=1}^{n}\log\frac{P_{i,1}(X_{i})}{P_{i,0}(X_{i})},$ where $X_{i}\sim P_{i,0}$ for all $i\in[n]$. Furthermore, we impose an assumption on the log-likelihood ratios.
@@
Assume that the log-likelihood ratio is bounded by some constants $L_{1},\dots,L_{n}\in\mathbb{R}_{>0}$ for each node $i$, that is,
@@
Assume that the consensus Markov chain is reversible.
Under assumption 2, the convergence of the power of $W$ is captured by $\rho\coloneqq\max\{\lambda_{2},\lvert\lambda_{n}\rvert\}$, where $\lambda_{k}$ is the $k$-th largest eigenvalue of $W$. Since the spectral gap, $1-\rho$, is related to the connectivity of the network, it is intuitive that the term emerges in the price of decentralization. However, we can bound the convergence of the power of $W$ with the network size $n$ even if the assumption is not guaranteed.
@@
Assume that each distribution of the log likelihood ratio, $\log\frac{P_{i,1}(X_{i})}{P_{i,0}(X_{i})}$, is non-lattice like for $i\in[n]$. By choosing $r=r^{*}$, the type-II error probability is upper bounded as
$\beta^{(t)*}_{\mathrm{cen}}(\epsilon)\leq\beta^{(t)*}_{i}(r^{*},\epsilon)\leq C^{(\mathrm{NP})}_{i}\cdot\beta^{(t)*}_{\mathrm{cen}}(\epsilon),$
where $\beta^{(t)*}_{\mathrm{cen}}(\epsilon)$ is the optimal type-II error probability in the centralized case and
$C^{(\mathrm{NP})}_{i}=\exp\left\{\frac{\rho}{1-\rho}\sqrt{\frac{1-\pi_{i}}{\pi_{i}}}\left(\sqrt{\sum_{j=1}^{n}\frac{1}{\pi_{j}}\left(D_{\mathrm{KL}}\left(P_{j,0}\|P_{j,1}\right)\right)^{2}}+\sqrt{\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}\right)+o(1)\right\}.$
The term, $\lambda_{k}$, is the $k$-th largest eigenvalue of the weight matrix $W$, and $\rho=\max\{\lambda_{2},\lvert\lambda_{n}\rvert\}$ is related to the connectivity of the network.
@@
For the constant punishment in Theorem 4, the term $\rho$ represents the connectivity of a graph such that a graph with higher connectivity (smaller $\rho$) obtains a smaller upper bound on the probability of error. For example, a complete graph with self-loops obtains the weight matrix $W=\frac{1}{n}\boldsymbol{1}\boldsymbol{1}^{\mathsf{T}}$ by making each node uniformly distributing weights to its neighbors. In this case, we have $\rho=0$ and thus the network obtains no constant punishment. Meanwhile, for a ring with each node giving both its neighbor half of its confidence, we have $\rho=\cos\frac{2\pi}{n}$ and the network suffers a larger price of decentralization as the network size $n$ grows.
Notice that the constant punishment may differ among the nodes with the stationary distribution at each node. Generally speaking, a node with a smaller corresponding $\pi_{i}$ gathers information slower since it gains less trust from the network compared to the others. For example, in Figure 2, a node that is far from the others in the network tends to obtain s smaller stationary distribution at it and suffer a larger price of decentralization.
Figure 2: A node with small $\pi_{i}$.
The reason the price of decentralization emerges as a constant term in the exponent of the error probability is also quite intuitive. Since we focus on the asymptotic analysis with respect to $t$ while the size of the network remains fixed, information needs only constant time to travel through the network regardless of the value of $t$. In fact, we can re-write Theorem 4 into the following form.
@@
$\beta^{(t)*}_{\mathrm{cen}}(\epsilon)\leq\beta^{(t)*}_{i}(r^{*},\epsilon)\overset{\cdot}{\leq}\beta^{(t-d_{i})*}_{\mathrm{cen}}(\epsilon),$
@@
$d_{i}=\frac{\log C^{(\mathrm{NP})}_{i}}{\sum_{j=1}^{n}D_{\mathrm{KL}}\left(P_{j,0}\|P_{j,1}\right)},$
is a constant with respect to $t$. The notation $a^{(t)}\overset{\cdot}{\leq}b^{(t)}$ means that $\lim_{t\rightarrow\infty}\frac{a^{(t)}}{b^{(t)}}\leq 1$.
@@
Assume that each distribution of the log likelihood ratio, $\log\frac{P_{i,1}(X_{i})}{P_{i,0}(X_{i})}$, is non-lattice like for $i\in[n]$. By choosing $r=r^{*}$, the Bayes risk is bounded as
$\mathsf{P}^{(t)*}_{\mathsf{e},\mathrm{cen}}(\xi)\leq\mathsf{P}^{(t)*}_{\mathsf{e},i}(r;\xi)\overset{\cdot}{\leq}C^{(\mathrm{B})}_{i}\cdot\mathsf{P}^{(t)*}_{\mathsf{e},\mathrm{cen}}(\xi)$
where $\mathsf{P}^{(t)*}_{\mathsf{e},\mathrm{cen}}(\xi)$ is the optimal Bayes risk in the centralized case and
$C^{(\mathrm{B})}_{i}=\exp\left\{\max\left\{\theta^{*},1-\theta^{*}\right\}\left(\frac{\rho}{1-\rho}\right)\sqrt{\frac{1-\pi_{i}}{\pi_{i}}\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}\right)}+o(1)\right\}$
$\theta^{*}=\underset{\theta\in[0,1]}{\arg\max}\ \left\{-\log\mathsf{E}_{X\sim P_{0}}\left[\left(\frac{P_{1}(X)}{P_{0}(X)}\right)^{\theta}\right]\right\}$
@@
In Theorem 4 and Theorem 5, the upper bounds hint that the effect of decentralization depends on the network connectivity with the term $\frac{\rho}{1-\rho}$. In the special case that $X_{i}^{(t)}$ follows Gaussian distributions for all $i\in[n]$, the optimal type-II error probability in the Neyman-Pearson problem is characterized, and it shows that the effect of decentralization relates to the network connectivity with the term $\frac{\rho^{2}}{1-\rho^{2}}$. When $\rho$ goes to zero, the term $\frac{\rho^{2}}{1-\rho^{2}}$ goes to zero faster than $\frac{\rho}{1-\rho}$, that is, in the Gaussian case, $\frac{\rho^{2}}{1-\rho^{2}}$ is a tighter characterization of the effect of decentralization. However, notice that in this case, our Assumption 3 is not satisfied. The result hints that our characterization fn the effect of decentralization is either incomplete due to the assumption or not tight enough.
@@
Assume that for each node $i$, the observations follow the i.i.d. Gaussian distributions, that is,
$\mathcal{H}_{0}:X^{(t)}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(-\mu,\sigma^{2})$
$\mathcal{H}_{1}:X^{(t)}_{i}\overset{i.i.d.}{\sim}\mathcal{N}(\mu,\sigma^{2}).$
@@
For Gaussian observations, the log-likelihood ratios follow another Gaussian distribution as well, and thus we characterize the optimal threshold with the inverse Q-function and plug the threshold in to directly calculate the corresponding type-II error probability. The rests are to approximate the Q-function and control the deviation of the Markov chain with some upper bounds. The detailed proof is provided in Appendix A-G. ∎

### v2

@@ Abstract
Index Terms:
Decentralized hypothesis testing, social learning, distributed learning, error exponent, higher-order asymptotics.
@@ I-A Contribution
In this work, the above questions are addressed in the case of binary detection. We propose a generalization of the social learning rule in [2] and characterize the error exponents using tools in large deviation theory [10]. As a result, the error exponents of the original learning rule in [2] are characterized, which turn out to be strictly smaller than the error exponents in the centralized case. The reason is that the decentralized sources are not weighted equally due to the convergence of the Markov chain governing the consensus. Figure 1 illustrates the gap in error exponents with a simple example. In the example, 300 scale-free networks with 100 nodes in each are sampled. Each node serves as an independent Bernoulli source having consensus weights uniformly distributed to its neighbors. Gathering the consensus weights into a right stochastic matrix, the Markov chain with such corresponding transition matrix induces a unique stationary distribution denoted by $\pi$ under some minor assumptions. The figure shows that the error exponent of the original learning rule decreases with the network imbalance. We quantify the imbalance of the network with the 2-norm between $\pi$ and the uniform stationary distribution with each entry being 0.01 for this case. Notice that only when the network is balanced, the original learning rule obtains the optimal error exponent depicted by the blue dashed line.
@@ I-C Paper Organization
The rest of this paper is organized as follows. In Section II, we formulate our problem and introduce the learning rule proposed in [2]. In Section III, we propose our modified learning rule and show our main results. The detailed proofs are provided in the appendix. We then propose alternative learning rules for estimating the needed parameters and discuss the convergence of the estimation in Section IV. In Section V, we provide simulation results on the impact of network imbalance, estimation, and quantization. In Section VI, we further discuss about various aspects of our results including removing the assumptions on the network and extension to the multiple hypothesis testing problems. Finally, we conclude the paper in Section VII.
@@ II-B Social Learning Rule
The coefficients $Z^{(t)}_{i,1},Z^{(t)}_{i,2}$ in steps 2) and 4) normalize the belief vectors such that they fall back into the $m$-dimensional probability simplex.
@@
Though [2] characterized the convergence performance of the belief vectors, they did not study the probability of error, which seems to be a more concerned perspective in the conventional hypothesis testing problem. As we briefly discussed in Section I, it can be observed from Figure 1 that if the network tends to be more imbalanced, the learning rule in [2] tends to suffer a larger gap in the error exponent compared to the centralized case, where a central node is assumed to be capable of gathering all the observations and likelihood functions to perform an optimal test. The gap between the error exponents motivates our modified learning rule introduced in the next section. It turns out that we can close the gap with slight modifications while keeping our learning rule working in a decentralized manner.
In the following, let us first introduce the log-belief ration test we consider in the rest of our work. For the centralized binary detection problem, the randomized likelihood ratio test is optimal (in the Neyman-Pearson problem and the Bayes setting). However, in the decentralized setting, none of the nodes knows the joint likelihood of all the observations in the network and thus no one can carry out the likelihood ratio test. Under the above-mentioned learning rule, we consider the binary hypothesis testing problem, and a natural test based on the private belief vector maintained by each node emerges, which is defined as follows.
@@ III-A Modified Learning Rule
Now that the additional parameters $r_{i}$’s emerge, the probability of error depends on the choice of the geometric weights. We formally define them as follows.
@@
$\mathsf{P}_{\mathsf{e},i}^{(t)*}(r;\xi)$
$\triangleq\min_{\eta_{i}^{(t)},\gamma_{i}^{(t)}}\left\{\xi_{0}\alpha_{i}^{(t)}(r;\eta_{i}^{(t)},\gamma_{i}^{(t)})+\xi_{1}\beta_{i}^{(t)}(r;\eta_{i}^{(t)},\gamma_{i}^{(t)})\right\}.$
@@ III-B First-Order Results
Suppose that Assumptions 1 and 2 hold. For the Neyman-Pearson problem, the type-II error exponent for each node $i$ is characterized as shown on the top of the next page.
$\forall\,\epsilon>0,\ \lim_{t\rightarrow\infty}-\frac{1}{t}\log\beta_{i}^{(t)*}(r,\epsilon)=\sup_{\lambda\geq 0}\left\{\sum_{j=1}^{n}\lambda\pi_{j}r_{j}\mathsf{E}_{P_{0}}\left[\log\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right]-\log\mathsf{E}_{P_{1}}\left[\left(\frac{P_{j,0}(X_{j})}{P_{j,1}(X_{j})}\right)^{\lambda\pi_{j}r_{j}}\right]\right\},\ i\in[n].$
@@
Suppose that Assumptions 1 and 2 hold. If each agent $i$ knows $\pi_{i}$, the value of the stationary distribution of the weight matrix at that node, by choosing $r=r^{*}$ where $r^{*}_{i}=c/\pi_{i}$ for all $i$ and some common constant $c\in\mathbb{R}$ among the nodes, we have
$\lim_{t\rightarrow\infty}-\frac{1}{t}\log\beta_{i}^{(t)*}(r^{*},\epsilon)$ $=\sum_{j=1}^{n}D_{\mathrm{KL}}\left(P_{j,0}\|P_{j,1}\right)$
$=D_{\mathrm{KL}}\left(P_{0}\|P_{1}\right)\quad\forall\,i\in[n].$
Here $P_{\theta}$ denotes the product distribution of $P_{1,\theta},...,P_{n,\theta}$ and $D_{\mathrm{KL}}$ denotes the Kullback–Leibler divergence.
@@
Suppose that Assumptions 1 and 2 hold. If each agent $i$ knows $\pi_{i}$, the value of the stationary distribution of the weight matrix at that node, by choosing $r=r^{*}$, for all prior $\xi$, we have
@@
where $\mathrm{CI}(P_{0},P_{1})$ denotes the Chernoff information between $P_{0}$ and $P_{1}$, that is,
$\mathrm{CI}(P_{0},P_{1})$
$=\max_{\lambda\in[0,1]}\left\{-\log\mathsf{E}_{P_{0}}\left[\left(\frac{P_{1}(\boldsymbol{X})}{P_{0}(\boldsymbol{X})}\right)^{\lambda}\right]\right\}$
$=\max_{\lambda\in[0,1]}\left\{-\log\mathsf{E}_{P_{0}}\left[\left(\frac{P_{1,1}(X_{1})\dots P_{n,1}(X_{n})}{P_{1,0}(X_{1})\dots P_{n,0}(X_{n})}\right)^{\lambda}\right]\right\}.$
@@ III-C Higher-Order Asymptotics
While Theorem 2 shows that decentralization does not affect the error exponent if we choose the geometric weights properly, further investigation on the probability of error may reveal the effect of decentralization. To characterize the higher-order asymptotics, we further impose the following assumption on the log-likelihood ratios.
@@
The log-likelihood ratio is bounded by some constants $L_{1},\dots,L_{n}>0$ for each node $i$, that is,
@@
The consensus Markov chain is reversible.
Under Assumption 4, the convergence of the power of $W$ is captured by $\rho\triangleq\max\{\lambda_{2},\lvert\lambda_{n}\rvert\}$, where $\lambda_{k}$ is the $k$-th largest eigenvalue of $W$, $k\in[n]$. Since the spectral gap, $1-\rho$, is related to the connectivity of the network, it is intuitive that the term emerges in the price of decentralization.
@@
Suppose that Assumptions 1, 2, 3, and 4 hold. Assume that each distribution of the log likelihood ratio, $\log\frac{P_{i,1}(X_{i})}{P_{i,0}(X_{i})}$, is non-lattice for $i\in[n]$. By choosing $r=r^{*}$, the type-II error probability is upper bounded as
$\beta^{(t)*}_{\mathrm{cen}}(\epsilon)\leq\beta^{(t)*}_{i}(r^{*},\epsilon)\leq C^{(\mathrm{NP})}_{i}\,\beta^{(t)*}_{\mathrm{cen}}(\epsilon),$
where $\beta^{(t)*}_{\mathrm{cen}}(\epsilon)$ is the optimal type-II error probability in the centralized case and the expression of the constant $C^{(\mathrm{NP})}_{i}$ is shown on the top of the next page.
@@
$C^{(\mathrm{NP})}_{i}=\exp\left\{\frac{\rho}{1-\rho}\sqrt{\frac{1-\pi_{i}}{\pi_{i}}}\left(\sqrt{\sum_{j=1}^{n}\frac{1}{\pi_{j}}\left(D_{\mathrm{KL}}\left(P_{j,0}\|P_{j,1}\right)\right)^{2}}+\sqrt{\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}}\right)+o(1)\right\},\quad i\in[n].$ (1)
For the constant penalty in Theorem 4, the term $\rho$ in (1) represents the connectivity of a graph such that a graph with higher connectivity (smaller $\rho$) obtains a smaller upper bound on the probability of error. For example, a complete graph with self-loops obtains the weight matrix $W=\frac{1}{n}\boldsymbol{1}\boldsymbol{1}^{\mathsf{T}}$ by making each node uniformly distributing weights to its neighbors. In this case, we have $\rho=0$ and thus the network obtains no constant penalty. Meanwhile, for a ring with each node giving both its neighbor half of its confidence, we have $\rho=\cos\frac{2\pi}{n}$ and the network suffers a larger price of decentralization as the network size $n$ grows.
Notice that the constant penalty may differ among the nodes with the stationary distribution at each node. Generally speaking, a node with a smaller corresponding $\pi_{i}$ gathers information slower since it gains less trust from the network compared to the others. For example, in Figure 2, a node that is far from the others in the network tends to obtain s smaller stationary distribution at it and suffer a larger price of decentralization.
Figure 2: A node (colored red) with small $\pi_{i}$.
The reason why the price of decentralization emerges as a constant term in the exponent of the error probability is also quite intuitive. Since we focus on the asymptotic analysis with respect to $t$ while the size of the network remains fixed, information needs only constant time to travel through the network regardless of the value of $t$. In fact, we can re-write Theorem 4 into the following form.
@@
$\beta^{(t)*}_{\mathrm{cen}}(\epsilon)\leq\beta^{(t)*}_{i}(r^{*},\epsilon)\lesssim\beta^{(t-d_{i})*}_{\mathrm{cen}}(\epsilon),$
@@
$d_{i}=\frac{\log C^{(\mathrm{NP})}_{i}}{\sum_{j=1}^{n}D_{\mathrm{KL}}\left(P_{j,0}\|P_{j,1}\right)},\ i\in[n],$
is a constant with respect to $t$. The notation $a^{(t)}\lesssim b^{(t)}$ means that $\lim_{t\rightarrow\infty}\frac{a^{(t)}}{b^{(t)}}\leq 1$.
@@
Suppose that Assumptions 1, 2, 3, and 4 hold. Assume that each distribution of the log likelihood ratio, $\log\frac{P_{i,1}(X_{i})}{P_{i,0}(X_{i})}$, is non-lattice for $i\in[n]$. By choosing $r=r^{*}$, the Bayes risk is bounded as
$\mathsf{P}^{(t)*}_{\mathsf{e},\mathrm{cen}}(\xi)\leq\mathsf{P}^{(t)*}_{\mathsf{e},i}(r;\xi)\lesssim C^{(\mathrm{B})}_{i}\,\mathsf{P}^{(t)*}_{\mathsf{e},\mathrm{cen}}(\xi)$
where $\mathsf{P}^{(t)*}_{\mathsf{e},\mathrm{cen}}(\xi)$ is the optimal Bayes risk in the centralized case and the expression of the constant $C^{(\mathrm{B})}_{i}$, $i\in[n]$, is shown below:
$\exp\left\{\max(\theta^{*},1-\theta^{*})\frac{\rho}{1-\rho}\sqrt{\frac{1-\pi_{i}}{\pi_{i}}\left(\sum_{j=1}^{n}\frac{1}{\pi_{j}}L_{j}^{2}\right)}+o(1)\right\},$
with $\theta^{*}=\underset{\theta\in[0,1]}{\arg\max}\ \left\{-\log\mathsf{E}_{X\sim P_{0}}\left[\left(\frac{P_{1}(X)}{P_{0}(X)}\right)^{\theta}\right]\right\}$.
@@
In Theorem 4 and Theorem 5, the upper bounds hint that the effect of decentralization depends on the network connectivity with the term $\frac{\rho}{1-\rho}$. In the special case that $X_{i}^{(t)}$ follows Gaussian distributions for all $i\in[n]$, the optimal type-II error probability in the Neyman-Pearson problem is characterized, and it shows that the effect of decentralization relates to the network connectivity with the term $\frac{\rho^{2}}{1-\rho^{2}}$. When $\rho$ goes to zero, the term $\frac{\rho^{2}}{1-\rho^{2}}$ goes to zero faster than $\frac{\rho}{1-\rho}$, that is, in the Gaussian case, $\frac{\rho^{2}}{1-\rho^{2}}$ is a tighter characterization of the effect of decentralization. However, notice that in this case, our Assumption 3 is not satisfied. The result hints that our characterization of the effect of decentralization is either incomplete due to the assumption or not tight enough.
@@
Suppose that Assumptions 1, 2, and 4 hold. Assume that for each node $i$, the observations follow the i.i.d. Gaussian distributions, that is,
[…]
@@
$\beta^{(t)*}_{i}(r^{*},\epsilon)\leq C^{(\mathrm{B,Gaussian})}_{i}\,\beta^{(t)*}_{\mathrm{cen}}(\epsilon)$
@@
For Gaussian observations, the log-likelihood ratios follow another Gaussian distribution as well, and thus we characterize the optimal threshold with the inverse Q-function and plug the threshold in to directly calculate the corresponding type-II error probability. The rest are to approximate the Q-function and control the deviation of the Markov chain with some upper bounds. The detailed proof is provided in Appendix A-F. ∎